Assume on January 1, 2014 you opened a fixed deposit account with a local bank and between January 1, 2014 and January 1, 2019 you made the following deposits and withdrawals:

Date

Deposits

Withdrawals

January 1, 2014

6,000

0

January 1, 2015

4,000

0

January 1, 2016

0

0

January 1, 2017

0

9,000

January 1, 2018

0

0

January 1, 2019

3,000

0

Assume an interest rate of r = 14% compounded annually.

(a) Write down the cash flow stream; (No need to key-in here.)

(b) What is the value in your account at the beginning of year 2020?

(Keep your answer to 2 decimal places, e.g. 987.12)
Answer #1

(a)

(b)

Interest Rate = 14 % compounded annually

Value of First Deposit on 1st Jan 2020 = 6000 x (1.14)^(2020 - 2014) = $ 13169.84

Value of Second Deposit on 1st Jan 2020 = 4000 x (1.14)^(2020-2015) = $ 7701.66

Value of First Withdrawal on 1st Jan 2020 = -9000 x (1.14)^(2020 - 2017) = - $ 13333.89

Value of Third Deposit on 1st Jan 2020 = 3000 x (1.14)^(2020-2019) = $ 3420

Net Account Value on 1st Jan 2020 = 13169.84 + 7701.66 - 13333.89 + 3420 = $ 10957.6

Final Value of the FIrst Deposit wor